What are the x-intercept(s) of the graph of the equation y − x2 = −6x − 7?

Answer by ewatrrr(24785)


Hi
What are the x-intercept(s) of the graph of the equation
y-x^2 = -6x-7
y = x^2-6x-7
Finding x-intercept when y = 0
x^2-6x-7 = 0
factoring
(x-7)(x+1) = 0
(x-7)=0 x = 7
(x+1)= 0 x= -1
x-intercept are(7,0) and (-1,0)